one of the acute angle of a right angle is 49 degrees. find the other acute angle
1
Expert's answer
2014-07-17T11:05:15-0400
In a triangle, the three interior angles always add to 180 degrees. If the given triangle is right, then one of angles is a right angle (that is a 90 degrees angle). So, we know angle of 49 degrees and that of 90 degrees.
Therefore, the other angle is 180-90-49=90-49=41 (degrees).
Answer: 41 degrees.